# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 6.1 Alternative descriptions
We start with a definition close to [Ber]. Let $\mathbf{Syn}^{\partial}_{\mathrm{an}}(\mathcal{E},\mathfrak{E},r)$ or simply $\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)$ be defined as the total complex associated to the morphism of complexes $(1-\frac{\Phi}{p^{r}},\gamma_{\rm an})$
@@
Assume that $(\mathcal{E},\nabla)$ has nilpotent residues (see Definition 1). Then, for every $i$ the morphisms $\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r)\to\overline{\mathbf{Syn}}(\mathcal{E},\mathfrak{E},r)$ and $\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)\to\overline{\mathbf{Syn}}(\mathcal{E},\mathfrak{E},r)$ are quasi-isomorphisms and, hence, induce isomorphisms on cohomology.
@@
The first statement follows as the natural map $\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r)\to\overline{\mathbf{Syn}}(\mathcal{E},\mathfrak{E},r)$ is a quasi-isomorphism. See the proof of [EY, Lemma 3.6] for the computation.
@@ 6.2 Relation with Hyodo-Kato cohomology with support
One defines complexes $\mathrm{R}\Gamma_{\mathrm{abs}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0},\mathfrak{E}^{\rm conv}(-\mathfrak{D})\bigr)$, $\mathrm{R}\Gamma_{\mathrm{abs}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0},\mathfrak{E}_{0}^{\rm conv}(-\mathfrak{D}_{0})\bigr)$ and $\mathrm{R}\Gamma_{\mathrm{HK}}\bigl(\mathcal{X}_{k}/\mathcal{S}_{0}^{0},\mathfrak{E}_{0}(-\mathfrak{D}_{0})\bigr)$ proceeding as in §5. We have the following :
@@ 6.3 Log-rigid cohomology with support conditions
First of all define the rigid syntomic cohomology of $U-\mathcal{D}_{k}\subset Y$ with compact support along $\mathcal{D}_{k}$ by $\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}}^{c\text{-}\infty}\left(\mathcal{E},\mathfrak{E},r\right):=\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}}\left(\mathcal{E}\left(-\mathcal{D}\right),\mathfrak{E}\left(-\mathfrak{D}\right),r\right)$.
@@
see §4.5. Define $\mathrm{dR}_{(U\subset Y,\alpha_{k})/K_{0}}\bigl(\mathfrak{E}^{\rm conv}(-\mathfrak{D})\bigr)_{\underline{\beta}}$ and $\mathrm{dR}_{(U\subset Y,\alpha_{k})/K}(\mathcal{E}(-\mathcal{D}))_{\underline{\beta}}$ to be the total complexes. Using Remark 4.1, one can prove that they compute the cohomology of the total complexes of the corresponding morphisms of logarithmic de Rham complexes of sheaves. Following [Ber, §3 & §4] denote by
@@
Consider the complex $\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r)$ introduced in Definition 5. By construction, we get a morphism of complexes
$\tau_{{\rm rig},c}\colon\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r)\longrightarrow\mathbf{Syn}^{c\text{-}\infty}(\mathcal{E},\mathfrak{E},r).$
@@
are quasi-isomorphisms. In particular, we have a quasi-isomorphism of complexes $\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r)\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathcal{E}^{\dagger},\mathfrak{E}^{\dagger},r).$
@@ 6.3.1 The case of $\mathcal{X}_{k}^{\rm sm}$
We now specialize to the case $U=\mathcal{X}_{k}^{\rm sm}$, $U^{o}=\mathcal{X}_{k}^{\rm sm,o}=\mathcal{X}_{k}^{\rm sm}\backslash\mathcal{D}_{k}$ and $Y=\mathcal{X}_{k}$. Let $W_{1},\ldots,W_{s}\subset\mathcal{X}_{k}$ be the irreducible components and let $W_{i}^{o}=W_{i}\cap U^{o}$. For each $i$ and each $\underline{\beta}$, replacing the tubes of $\mathcal{X}_{\underline{\beta},k}$ and of $Z_{k}$ in the rigid analytic fiber of $\mathfrak{X}_{\underline{\beta}}$, resp. of $\mathcal{X}$ with the tubes of $W_{i}\cap\mathcal{X}_{\underline{\beta},k}$ and of $Z_{k}\cap W_{i}\cap\mathcal{X}_{\underline{\beta},k}$ in the construction above we get complexes $\mathrm{R\Gamma}_{\rm rig,c}\bigl((W_{i}^{o}\subset W_{i},\alpha_{k})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$, $\mathrm{R\Gamma}_{\rm rig,c}\bigl((W_{i}^{o}\subset W_{i},\alpha_{k})/K,\mathcal{E}\bigr)$ and $\mathbf{Syn}_{{\rm rig},W_{i}^{o}\subset W_{i},\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r)$. We also have morphisms
@@
and, hence, $\xi\colon\mathbf{Syn}_{{\rm rig},\mathcal{X}_{k}^{\rm sm,o}\subset\mathcal{X}_{k},\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r)\longrightarrow\oplus_{i=1}^{s}\mathbf{Syn}_{{\rm rig},W_{i}^{o}\subset W_{i},\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r)$.
@@
$\tau_{\mathrm{rig},U,\ell,c}^{i}\colon\mathrm{H}^{i}\left(\mathbf{Syn}_{\mathrm{rig},U^{o}\subset W_{\ell},\alpha_{k}}^{c}\left(\mathcal{E},\mathfrak{E},r\right)\right)\longrightarrow\mathrm{H}^{i}\left(\mathbf{Syn}^{c\text{-}\infty}\left(\mathcal{E},\mathfrak{E},r\right)\right).$ (1)
@@ 6.4 Coefficients with good reduction
Define complexes $\mathrm{R}\Gamma_{\mathrm{rig}}\bigl((U\subset Y,\alpha_{k}^{\rm hor})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$ and $\mathbf{Syn}_{{\rm rig},U\subset Y,\alpha_{k}^{\rm hor}}\bigl(\mathcal{E},\mathfrak{E},r\bigr)$ proceeding as in §4.5 and $\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}^{hor}}^{c\text{-}\infty}\left(\mathcal{E},\mathfrak{E},r\right):=\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}^{hor}}\left(\mathcal{E}\left(-\mathcal{D}\right),\mathfrak{E}\left(-\mathfrak{D}\right),r\right)$, $\mathrm{R}\Gamma_{\mathrm{rig},c}\bigl((U^{o}\subset Y,\alpha_{k}^{\rm hor})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$ and $\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}\bigl(\mathcal{E},\mathfrak{E},r\bigr)$ as in §6.3. In terms of the complexes of loc. cit. we take the logarithmic complexes only with respect to the variables $Y_{1},\ldots,Y_{b}$ and not with respect to the variables $X_{1},\ldots,X_{a}$. In particular, we have natural morphisms of complexes $\Xi$ from
@@ 6.5 An acyclicity result
(iv) Suppose that $0\longrightarrow M^{\prime}\longrightarrow M\longrightarrow M^{\prime\prime}\longrightarrow 0$ is an exact sequence of $R$-modules and that $\varphi_{i}\left(M^{\prime}\right)\subset M^{\prime}$ for every $i=1,\ldots,n$. Write $\left\{\varphi_{1}^{\prime},\ldots,\varphi_{n}^{\prime}\right\}\subset\mathrm{End}_{R}\left(M^{\prime}\right)$ and $\left\{\varphi_{1}^{\prime\prime},\ldots,\varphi_{n}^{\prime\prime}\right\}\subset\mathrm{End}_{R}\left(M^{\prime\prime}\right)$ for the induced morphisms. Then we have an exact sequence of complexes of $R$-modules
@@ 7 Period sheaves
Following [DLLZb, §2.2] we have period sheaves $\widehat{\mathcal{O}}_{X}$ and $\mathbb{A}_{\mathrm{inf}}$ on $X_{\mathrm{pket}}$ with a morphism $\vartheta\colon\mathbb{A}_{\mathrm{inf}}\to\widehat{\mathcal{O}}_{X}^{+}$ with locally principal kernel. From these we get filtered period sheaves $\mathbb{B}_{\mathrm{crys}}$ and $\mathbb{B}_{\mathrm{dR}}$. For every $J$ we have the sheaves $\widehat{\mathcal{O}}_{X_{J}}^{\partial,+}$, $\mathbb{A}_{\mathrm{inf},J}^{\partial}$ with the morphism $\vartheta_{J}^{\partial}\colon\mathbb{A}_{\mathrm{inf},J}^{\partial}\to\widehat{\mathcal{O}}_{X_{J}}^{\partial,+}$ and the filtered period sheaves $\mathbb{B}_{\mathrm{crys},J}^{\partial}$ and $\mathbb{B}_{\mathrm{dR},J}^{\partial}$ on $X_{\mathrm{pket}}$ obtained as push-forward via $\iota_{J}\colon X_{J}^{\partial}\to X$ of the corresponding period sheaves on $X^{\partial}_{J,\mathrm{pket}}$ ; see [LLZ, Def. 2.5.2] for the de Rham version. For the crystalline version one defines $\mathbb{A}_{\mathrm{crys},J}^{\partial}$ as in [TT, Def. 2.1] as the $p$-adic completion of the DP envelope of the ideal $\mathrm{ker}(\vartheta_{J}^{\partial})$, which is locally principal by [LLZ, Cor. 2.5.3]. Then $\mathbb{B}_{\mathrm{crys},J}^{\partial}=\mathbb{A}_{\mathrm{crys},J}^{\partial}[t^{-1}]$, with $t$ Fontaine’s element. The sheaves $\widehat{\mathcal{O}}_{X}$, $\mathbb{B}_{\mathrm{crys}}$, $\mathbb{B}_{\mathrm{dR}}$ etc. correspond to the case $J=\emptyset$. Set $\mathbb{B}_{\mathrm{crys},(a)}^{\partial}:=\oplus_{J\subset I,|J|=a}\mathbb{B}_{\mathrm{crys},J}^{\partial}$ and $\mathbb{B}_{\mathrm{dR},(a)}^{\partial}=\oplus_{J\subset I,|J|=a}\mathbb{B}_{\mathrm{dR},J}^{\partial}$. Define the period sheaves with compact support
$\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}:=\mathbb{B}_{\mathrm{crys},(\bullet)}^{\partial},\qquad\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}:=\mathbb{B}_{\mathrm{dR},(\bullet)}^{\partial}$
with filtration $\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}:=\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{crys},(\bullet)}^{\partial}$ and $\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}:=\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{dR},(\bullet)}^{\partial}$. The crystalline period sheaves are endowed with Frobenius endomorphisms and we have injections, strictly compatible with filtrations,
$\mathbb{B}_{\mathrm{crys}}\subset\mathbb{B}_{\mathrm{dR}},\qquad\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}\subset\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}.$
@@
In [LLZ, Def. 3.3.2] one finds the definition of $\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}$ as the kernel of $\mathbb{B}_{\mathrm{dR},(0)}^{\partial}\to\mathbb{B}_{\mathrm{dR},(1)}^{\partial}$. It is proven in [LLZ, Lemma 3.3.7] that the natural map from this to $\mathbb{B}_{\mathrm{dR},(\bullet)}^{\partial}$ is an isomorphism.
In [DLLZb] and [LLZ] relative versions of these sheaves are introduced, $\mathcal{O}\mathbb{B}_{\mathrm{dR}}$ and $\mathcal{O}\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}$. Furthermore by [DLLZb, Cor. 2.4.2] and [LLZ, Cor. 2.5.13] we have quasi-siomorphisms of filtered complexes
$\mathbb{B}_{\mathrm{dR}}\longrightarrow\mathcal{O}\mathbb{B}_{\mathrm{dR}}\otimes_{\mathcal{O}_{X}}\Omega^{\log,\bullet}_{X},\qquad\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}\longrightarrow\mathcal{O}\mathbb{B}_{\mathrm{dR}}^{c\text{-}\infty}\otimes_{\mathcal{O}_{X}}\Omega^{\log,\bullet}_{X}$.
Under Assumption 1.1 and fixing $\beta\in B$ one has also relative versions ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}$ and ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}$ on $X_{\beta,\mathrm{pket}}$ as follows. Let $w\colon X_{\beta}\to\mathcal{X}_{\beta,k}$ be the specialization map. It induces a mophism on étale sites. Define ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}$ as the $p$-adic completion of the log DP envelope of
@@
see [AI, Lemma 2.16]. They are endowed with Frobenius morphisms induced by the given Frobenius on $\mathfrak{X}_{\beta}$ and Frobenius on $\mathbb{A}_{\mathrm{inf}}$. Define ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}$ and ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys},J}^{c\text{-}\infty}$ as ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}[t^{-1}]$ and ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys},J}^{\partial}[t^{-1}]$ respetively. We consider the de Rham complexes
${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\otimes_{\mathcal{O}_{\mathfrak{X}_{\beta}}}\Omega_{\mathfrak{X}_{\beta}}^{\log,\bullet},\qquad{\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}^{c\text{-}\infty}\otimes_{\mathcal{O}_{\mathfrak{X}_{\beta}}}\Omega_{\mathfrak{X}_{\beta}}^{\log,\bullet}$
@@
sending $Z\mapsto\pi$, in the sense of [AI, Lemma 2.16]. Let $B_{\rm log}$ be $A_{\rm log}[t^{-1}]$. As explained in [AI, §2.1.1] this coincides with the ring $\widehat{B}_{\mathrm{st}}$ introduced by Breuil in [Breuil]. As explained in loc. cit., it contains Fontaine’s classical period ring $B_{\rm st}$ as the subring of $\widehat{B}_{\mathrm{st}}$ on which the monodromy operator $N$ is nilpotent. One can use it to characterize semistable $p$-adic representations of the absolute Galois group of $K$.

### v2

@@ 6.1 Alternative descriptions
We start with a definition close to [BER86]. Let $\mathbf{Syn}^{\partial}_{\mathrm{an}}(\mathcal{E},\mathfrak{E},r)$ or simply $\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)$ be defined as the total complex associated to the morphism of complexes $(1-\frac{\Phi}{p^{r}},\gamma_{\rm an})$
@@
Assume that $(\mathcal{E},\nabla)$ has nilpotent residues (see Definition 1). Then, for every $i$ the morphisms $\mathbf{Syn}^{c-\infty}(\mathcal{E},\mathfrak{E},r)\to\overline{\mathbf{Syn}}(\mathcal{E},\mathfrak{E},r)$ and $\mathbf{Syn}^{\partial}(\mathcal{E},\mathfrak{E},r)\to\overline{\mathbf{Syn}}(\mathcal{E},\mathfrak{E},r)$ are quasi-isomorphisms and, hence, induce isomorphisms on cohomology.
@@
The first statement follows as the natural map $\mathbf{Syn}^{c-\infty}(\mathcal{E},\mathfrak{E},r)\to\overline{\mathbf{Syn}}(\mathcal{E},\mathfrak{E},r)$ is a quasi-isomorphism. See the proof of [EY20, Lemma 3.6] for the computation.
@@ 6.2 Relation with Hyodo-Kato cohomology with support
One defines complexes $\mathrm{R}\Gamma_{\mathrm{abs}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0},\mathfrak{E}^{\rm conv}(-\mathfrak{D})\bigr)$, $\mathrm{R}\Gamma_{\mathrm{abs}}\bigl((\mathcal{X}_{k},\alpha_{\mathcal{X}_{k}})/\mathcal{S}_{0},\mathfrak{E}_{0}(-\mathfrak{D}_{0})\bigr)$ and $\mathrm{R}\Gamma_{\mathrm{HK}}\bigl(\mathcal{X}_{k}/\mathcal{S}_{0}^{0},\mathfrak{E}_{0}(-\mathfrak{D}_{0})\bigr)$ proceeding as in §5. We have the following :
@@ 6.3 Log-rigid cohomology with support conditions
First of all we introduce a rigid syntomic cohomology of $U-\mathcal{D}_{k}\subset Y$ with compact support along $\mathcal{D}_{k}$ by setting $\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}}^{c-\infty}\left(\mathcal{E},\mathfrak{E},r\right):=\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}}\left(\mathcal{E}\left(-\mathcal{D}\right),\mathfrak{E}\left(-\mathfrak{D}\right),r\right)$.
@@
see §4.5. Define
$\mathrm{dR}_{(U\subset Y,\alpha_{k})/K_{0}}\bigl(\mathfrak{E}^{\rm conv}(-\mathfrak{D})\bigr)_{\underline{\beta}},\quad\mathrm{dR}_{(U\subset Y,\alpha_{k})/K}(\mathcal{E}(-\mathcal{D}))_{\underline{\beta}}$
to be the total complexes. Using Remark 4.1, one can prove that they compute the cohomology of the total complexes of the corresponding morphisms of logarithmic de Rham complexes of sheaves. Following [BER86, §3 & §4] denote by
@@
Consider the complex $\mathbf{Syn}^{c-\infty}(\mathcal{E},\mathfrak{E},r)$ introduced in Definition 5. By construction, we get a morphism of complexes
$\tau_{{\rm rig},c}\colon\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r)\longrightarrow\mathbf{Syn}^{c-\infty}(\mathcal{E},\mathfrak{E},r).$
@@
are quasi-isomorphisms. In particular, we have a natural quasi-isomorphism of complexes
$\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r)\longrightarrow\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}^{c}(\mathcal{E}^{\dagger},\mathfrak{E}^{\dagger},r).$
@@ 6.3.1 The case of $\mathcal{X}_{k}^{\rm sm}$
We now specialize to the case $U=\mathcal{X}_{k}^{\rm sm}$, $U^{o}=\mathcal{X}_{k}^{\rm sm,o}=\mathcal{X}_{k}^{\rm sm}\backslash\mathcal{D}_{k}$ and $Y=\mathcal{X}_{k}$. Let $W_{1},\ldots,W_{s}\subset\mathcal{X}_{k}$ be the irreducible components and let $W_{i}^{o}=W_{i}\cap U^{o}$. For each $i$ and each $\underline{\beta}$, replacing the tubes of $\mathcal{X}_{\underline{\beta},k}$ and of $Z_{k}$ in the rigid analytic fiber of $\mathfrak{X}_{\underline{\beta}}$, resp. of $\mathcal{X}$ with the tubes of $W_{i}\cap\mathcal{X}_{\underline{\beta},k}$ and of $Z_{k}\cap W_{i}\cap\mathcal{X}_{\underline{\beta},k}$ in the construction above we get complexes $\mathrm{R\Gamma}_{\rm rig,c}\bigl((W_{i}^{o}\subset W_{i},\alpha_{k})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$, $\mathrm{R\Gamma}_{\rm rig,c}\bigl((W_{i}^{o}\subset W_{i},\alpha_{k})/K,\mathcal{E}\bigr)$ and $\mathbf{Syn}_{{\rm rig},W_{i}^{o}\subset W_{i},\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r)$. We also have natural morphisms from
@@
and, hence,
$\xi\colon\mathbf{Syn}_{{\rm rig},\mathcal{X}_{k}^{\rm sm,o}\subset\mathcal{X}_{k},\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r)\longrightarrow\oplus_{i=1}^{s}\mathbf{Syn}_{{\rm rig},W_{i}^{o}\subset W_{i},\alpha_{k}}^{c}(\mathcal{E},\mathfrak{E},r).$
@@
$\tau_{\mathrm{rig},U,\ell,c}^{i}\colon\mathrm{H}^{i}\left(\mathbf{Syn}_{\mathrm{rig},U^{o}\subset W_{\ell},\alpha_{k}}^{c}\left(\mathcal{E},\mathfrak{E},r\right)\right)\longrightarrow\mathrm{H}^{i}\left(\mathbf{Syn}^{c-\infty}\left(\mathcal{E},\mathfrak{E},r\right)\right).$ (1)
@@ 6.4 Coefficients with good reduction
We define complexes $\mathrm{R}\Gamma_{\mathrm{rig}}\bigl((U\subset Y,\alpha_{k}^{\rm hor})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$ and $\mathbf{Syn}_{{\rm rig},U\subset Y,\alpha_{k}^{\rm hor}}\bigl(\mathcal{E},\mathfrak{E},r\bigr)$ proceeding as in §4.5 and $\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}^{hor}}^{c\text{-}\infty}\left(\mathcal{E},\mathfrak{E},r\right):=\mathbf{Syn}_{\mathrm{rig},U\subset Y,\alpha_{k}^{hor}}\left(\mathcal{E}\left(\mathcal{D}\right),\mathfrak{E}\left(\mathfrak{D}\right),r\right)$, $\mathrm{R}\Gamma_{\mathrm{rig},c}\bigl((U^{o}\subset Y,\alpha_{k}^{\rm hor})/K_{0},\mathfrak{E}^{\rm conv}\bigr)$ and $\mathbf{Syn}_{{\rm rig},U^{o}\subset Y,\alpha_{k}}\bigl(\mathcal{E},\mathfrak{E},r\bigr)$ as in §6.3. In terms of the complexes of loc. cit. we take the logarithmic complexes only with respect to the variables $Y_{1},\ldots,Y_{b}$ and not with respect to the variables $X_{1},\ldots,X_{a}$. In particular, we have natural morphisms of complexes $\Xi$ from
@@ 6.5 An acyclicity result
(iv) Suppose that
$0\longrightarrow M^{\prime}\longrightarrow M\longrightarrow M^{\prime\prime}\longrightarrow 0$
is an exact sequence of $R$-modules and that $\varphi_{i}\left(M^{\prime}\right)\subset M^{\prime}$ for every $i=1,\ldots,n$. Write $\left\{\varphi_{1}^{\prime},\ldots,\varphi_{n}^{\prime}\right\}\subset\mathrm{End}_{R}\left(M^{\prime}\right)$ and $\left\{\varphi_{1}^{\prime\prime},\ldots,\varphi_{n}^{\prime\prime}\right\}\subset\mathrm{End}_{R}\left(M^{\prime\prime}\right)$ for the induced morphisms. Then we have an exact sequence of complexes of $R$-modules
@@ 7 Period sheaves
Following [DLL+23, §2.2] we have period sheaves $\widehat{\mathcal{O}}_{X}$ and $\mathbb{A}_{\mathrm{inf}}$ on $X_{\mathrm{pket}}$ with a morphism $\vartheta\colon\mathbb{A}_{\mathrm{inf}}\to\widehat{\mathcal{O}}_{X}^{+}$ with locally principal kernel. From these we get filtered period sheaves $\mathbb{B}_{\mathrm{crys}}$ and $\mathbb{B}_{\mathrm{dR}}$. For every $J$ we have the sheaves $\widehat{\mathcal{O}}_{X_{J}}^{\partial,+}$, $\mathbb{A}_{\mathrm{inf},J}^{\partial}$ with the morphism $\vartheta_{J}^{\partial}\colon\mathbb{A}_{\mathrm{inf},J}^{\partial}\to\widehat{\mathcal{O}}_{X_{J}}^{\partial,+}$ and the filtered period sheaves $\mathbb{B}_{\mathrm{crys},J}^{\partial}$ and $\mathbb{B}_{\mathrm{dR},J}^{\partial}$ on $X_{\mathrm{pket}}$ obtained as push-forward via $\iota_{J}\colon X_{J}^{\partial}\to X$ of the corresponding period sheaves on $X^{\partial}_{J,\mathrm{pket}}$ ; see [LLZ23, Def. 2.5.2] for the de Rham version. For the crystalline version one defines $\mathbb{A}_{\mathrm{crys},J}^{\partial}$ as in [TT19, Def. 2.1] as the $p$-adic completion of the DP envelope of the ideal $\mathrm{ker}(\vartheta_{J}^{\partial})$, which is locally principal by [LLZ23, Cor. 2.5.3]. Then $\mathbb{B}_{\mathrm{crys},J}^{\partial}=\mathbb{A}_{\mathrm{crys},J}^{\partial}[t^{-1}]$, with $t$ Fontaine’s element. The sheaves $\widehat{\mathcal{O}}_{X}$, $\mathbb{B}_{\mathrm{crys}}$, $\mathbb{B}_{\mathrm{dR}}$ etc. correspond to the case $J=\emptyset$. Set $\mathbb{B}_{\mathrm{crys},(a)}^{\partial}:=\oplus_{J\subset I,|J|=a}\mathbb{B}_{\mathrm{crys},J}^{\partial}$ and $\mathbb{B}_{\mathrm{dR},(a)}^{\partial}=\oplus_{J\subset I,|J|=a}\mathbb{B}_{\mathrm{dR},J}^{\partial}$. Define the period sheaves with compact support
$\mathbb{B}_{\mathrm{crys}}^{c-\infty}:=\mathbb{B}_{\mathrm{crys},(\bullet)}^{\partial},\qquad\mathbb{B}_{\mathrm{dR}}^{c-\infty}:=\mathbb{B}_{\mathrm{dR},(\bullet)}^{\partial}$
with filtration $\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{crys}}^{c-\infty}:=\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{crys},(\bullet)}^{\partial}$ and $\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{dR}}^{c-\infty}:=\mathrm{Fil}^{r}\mathbb{B}_{\mathrm{dR},(\bullet)}^{\partial}$. The crystalline period sheaves are endowed with Frobenius endomorphisms and we have injections, strictly compatible with filtrations,
$\mathbb{B}_{\mathrm{crys}}\subset\mathbb{B}_{\mathrm{dR}},\qquad\mathbb{B}_{\mathrm{crys}}^{c-\infty}\subset\mathbb{B}_{\mathrm{dR}}^{c-\infty}.$
@@
In [LLZ23, Def. 3.3.2] one finds the definition of $\mathbb{B}_{\mathrm{dR}}^{c-\infty}$ as the kernel of $\mathbb{B}_{\mathrm{dR},(0)}^{\partial}\to\mathbb{B}_{\mathrm{dR},(1)}^{\partial}$. It is proven in [LLZ23, Lemma 3.3.7] that the natural map from this to $\mathbb{B}_{\mathrm{dR},(\bullet)}^{\partial}$ is an isomorphism.
In [DLL+23] and [LLZ23] relative versions of these sheaves are introduced, $\mathcal{O}\mathbb{B}_{\mathrm{dR}}$ and $\mathcal{O}\mathbb{B}_{\mathrm{dR}}^{c-\infty}$. Furthermore by [DLL+23, Cor. 2.4.2] and [LLZ23, Cor. 2.5.13] we have quasi-siomorphisms of filtered complexes
$\mathbb{B}_{\mathrm{dR}}\longrightarrow\mathcal{O}\mathbb{B}_{\mathrm{dR}}\otimes_{\mathcal{O}_{X}}\Omega^{\log,\bullet}_{X},\qquad\mathbb{B}_{\mathrm{dR}}^{c-\infty}\longrightarrow\mathcal{O}\mathbb{B}_{\mathrm{dR}}^{c-\infty}\otimes_{\mathcal{O}_{X}}\Omega^{\log,\bullet}_{X}$.
Under Assumption 1.1 and fixing $\beta\in B$ one has also relative versions ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}$ and ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}^{c-\infty}$ on $X_{\beta,\mathrm{pket}}$ as follows. Let $w\colon X_{\beta}\to\mathcal{X}_{\beta,k}$ be the specialization map. It induces a mophism on étale sites. Define ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}$ as the $p$-adic completion of the log DP envelope of
@@
see [AI12, Lemma 2.16]. They are endowed with Frobenius morphisms induced by the given Frobenius on $\mathfrak{X}_{\beta}$ and Frobenius on $\mathbb{A}_{\mathrm{inf}}$. Define ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}$ and ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys},J}^{c-\infty}$ as ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys}}[t^{-1}]$ and ${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{A}_{\mathrm{crys},J}^{\partial}[t^{-1}]$ respetively. We consider the de Rham complexes
${\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}\otimes_{\mathcal{O}_{\mathfrak{X}_{\beta}}}\Omega_{\mathfrak{X}_{\beta}}^{\log,\bullet},\qquad{\mathcal{O}}_{\mathfrak{X}_{\beta}}\mathbb{B}_{\mathrm{crys}}^{c-\infty}\otimes_{\mathcal{O}_{\mathfrak{X}_{\beta}}}\Omega_{\mathfrak{X}_{\beta}}^{\log,\bullet}$
@@
sending $Z\mapsto\pi$, in the sense of [AI12, Lemma 2.16]. Let $B_{\rm log}$ be $A_{\rm log}[t^{-1}]$. As explained in [AI12, §2.1.1] this coincides with the ring $\widehat{B}_{\mathrm{st}}$ introduced by Breuil in [BRE97]. As explained in loc. cit., it contains Fontaine’s classical period ring $B_{\rm st}$ as the subring of $\widehat{B}_{\mathrm{st}}$ on which the monodromy operator $N$ is nilpotent. One can use it to characterize semistable $p$-adic representations of the absolute Galois group of $K$.